OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. If p is the number of pounds of chocolate-covered pretzels Troy purchases, what expression could be used to represent the number of pounds of gumdrops he purchases? (Points : 1) p + 1.8 p –1.8 1.8 – p 1.8p

OpenStudy (anonymous):

p-1.8 is your answer

OpenStudy (anonymous):

thanks i have a few more can u help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. If p = pounds of jelly beans, which equation in simplest form represents the situation described. (Points : 1) 0.31p + 1.41 = 1.56 0.31p + 1.41 = 2.1 –0.31p + 2.06 = 1.56 –0.31p + 2.06 = 2.1

OpenStudy (anonymous):

0.31p + 1.41 = 1.56 0.31p + 1.41 = 2.1 –0.31p + 2.06 = 1.56 –0.31p + 2.06 = 2.1

OpenStudy (anonymous):

im cinfuse why it does the question marks

OpenStudy (anonymous):

are you in k12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

9th wbu

OpenStudy (anonymous):

8th can you answer my question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

u have skype?

OpenStudy (anonymous):

why

OpenStudy (anonymous):

i can't solve this one i am stumped

OpenStudy (anonymous):

i like skype and if u ever need help u can ask me anytime :D and whats the question

OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 3x + 2 • x • 11 = 25x. 1. 3x + 2 • x • 11 = 3x + 2 • 11 • x 1. ? 2. 3x + 2 • x • 11 = 3x + 22x 2. Multiplication 3. 3x + 2 • x • 11 = 25x 3. ? A. Commutative Property of Addition, Addition B. Commutative Property of Multiplication, Addition C. Associative Property of Multiplication, Addition D. Commutative Property of Multiplication, Multiplication

OpenStudy (anonymous):

just go to the one i posted

OpenStudy (anonymous):

go to it and mention me